OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (gavlois1):

For functions of the format: \[y = mx+b\] m is equal to the slope of the line while b is actually the y-intercept. So right from the start, we can tell that from: \[y = -2x-3\] -2 is the slope (if required) and -3 is the y-intercept. To solve for the x,y coordinates of the y variable to -3 to find its x counterpart: \[y = -2x-3\] \[-3 = -2x -3\] We move the -3 on the right side to the left: \[-3 + 3 = -2x\] \[0 = -2x\] Divide both sides by -2: \[\frac{ 0 }{ -2 } = \frac{ -2x }{ -2 }\] Therefore, the x-coordinate of the y-intercept is \[0 = x\] Then we can safely say the coordinates of the y-intercept is (0,-3). However, the easiest way to do this is that we know intercepts are always on the axis, therefore the other coordinate is 0. y-intercepts have an x coordinate of 0 and x-intercepts have a y coordinate of 0. Using that idea, we can solve for the x -intercept. We set y to 0 since the intercept calls for the point of intersection at the x-axis. \[0 = -2x -3\] \[3 = -2x\] \[-\frac{ 3 }{ 2 }=x\] And so, we can say that the coordinates of the x-intercept is (-3/2 , 0)